# Style-Friendly SNR Sampler for Style-Driven Generation

### v2

@@ 4.1 Analysis of Style-Friendly SNR Sampler
In Fig. 5 and Fig. 6, we conduct experiments to analyze the impact of varying the parameters of our Style-friendly SNR sampler—specifically, the mean ($\mu$) and the standard deviation ($\sigma$) of the log-SNR sampling distribution, as well as the LoRA rank. We refer to the noise level distribution used for the SD3 pre-training as the SD3 sampler [8].
@@
The mean $\mu$ is the most critical design choice influencing style learning. We experiment with $\mu$ values ranging from 0 to $-8$ for both FLUX-dev and SD3.5-8B. As shown in Fig. 5, increasing $\mu$ leads the models to increasingly fail to learn reference styles. In Fig. 6(a), the DINO similarity decreases with increasing $\mu$, indicating inferior style alignment.
Conversely, in Fig. 5, when $\mu$ is set to $-6$ or lower (more negative), the models begin to capture and reflect the reference styles effectively. In Fig. S5, we also show the effect of $\mu$ on object references.
@@
We also investigate the effect of varying the standard deviation $\sigma$ of the log-SNR sampling distribution in Fig. 6(b). When $\sigma$ is less than 2, the model learns from a narrower interval, which can reduce style alignment due to less diversity in noise levels. This suggests that while $\sigma$ influences the breadth of noise levels, maintaining a moderate $\sigma$ (e.g., $\sigma=2$) provides a good balance between focusing on critical noise levels and learning on sufficiently diverse noise levels for style learning.
@@ 4.2 Qualitative Results
In Fig. 7, our Style-friendly SNR sampler accurately captures the styles of reference images, reflecting stylistic features including color schemes, layouts, illumination, and brushstrokes. In contrast, fine-tuning FLUX-dev with the standard SD3 sampler often fails to capture key stylistic components, such as layouts (columns 1, 2) and color schemes (columns 3-7). [^1] Note that some open-source implementations (e.g., Hugging Face Diffusers [58] v 0.31) may omit timestep shifting during fine-tuning. Readers should be aware of this potential omission when reproducing results.
Fine-tuning FLUX-dev with DCO struggles to learn the reference styles due to strong regularization that prevents significant deviation from the pre-trained model. IP-Adapter with FLUX-dev and RB-Modulation rely on embeddings of CLIP [39] and CSD [52], which may not capture fine stylistic details, leading to less accurate style reproduction. Style-Aligned shares self-attention features within the diffusion model, which can cause artifacts such as destroyed structure (columns 1, 2) or duplicated limbs (columns 4-7) when attention features conflict.
Figure 7: Qualitative comparison. All samples are generated with the same seed. Please zoom in.
@@
We further conduct a user study to quantify human preferences using Amazon Mechanical Turk. Following previous work [51], we compare our method to each method with two separate questionnaires. According to the reference style image and target text prompt, users are asked to select which of the two generated images is more similar to the style in the reference image and represents the target text prompt. We obtain 450 answers from 150 participants for each comparison, and the results are presented in Tab. 1. Our method outperforms prior arts in both aspects ($p<0.05$ in the Wilcoxon signed-rank test), consistent with the qualitative results and demonstrates the superiority of our method in learning stylistic elements. More details on our user study are provided in Sec. C.3.
4.3 Quantitative Results
In Tab. 2, we evaluate our method and prior works using DINO [2] and CLIP image similarities (CLIP-I) to assess style alignment, and CLIP text-image similarity (CLIP-T) for text alignment. Our method achieves the highest DINO and CLIP-I scores across all backbones, demonstrating its superior ability to capture styles from the reference images.
While our CLIP-T score is slightly lower compared to some methods, we already showed superior text alignment in human evaluation (Tab. 1). This implies that methods that fail to capture styles often generate the most common interpretation of the prompt, leading to higher CLIP-T scores due to bias towards typical representations. Overall, our quantitative results confirm that our method accurately reflects both styles and texts.
@@ 4.3 Quantitative Results
SD3 sampler [8] SD3.5 0.424 0.670 0.350
Style-friendly SD3.5 0.489 0.698 0.349
IP-Adapter [61] FLUX-dev 0.361 0.656 0.354
DCO [30] FLUX-dev 0.373 0.643 0.353
SD3 sampler [8] FLUX-dev 0.373 0.645 0.350
Style-friendly FLUX-dev 0.461 0.686 0.344
Table 2: Quantitative comparison. Style alignment (DINO and CLIP-I) and text alignment (CLIP-T) with 18 styles from [51]. Our style-friendly exhibits superior style-alignment scores.
@@ 4.4 Applications
Multi-Panel. Dreambooth [45] shows a multi-panel comic by generating each panel with a model fine-tuned on a cartoon character. However, in the first row of Fig. 8, we treat these multiple panels as a single image during fine-tuning. When a new subject is specified with a target prompt, we generate the subject across multiple comic-style panels simultaneously. This approach allows for the generation of coherent multi-panel comics using a single reference.
Typography. Our method can also be extended to typography, including meme generation as shown in Fig. 1. Leveraging the spelling capabilities of recent models [29, 57], we generated customized typography with unique styles, and the results are showcased in the second row of Fig. 8. This versatility enables users to effortlessly generate a wide range of customized textual elements.
Figure 8: Multi-panel and typography. First row demonstrates generating multiple coherent images as a single image. Second row shows customized typography with a unique style.
@@ 5.2 Style-Driven Generation
With advancements in text-to-image models, practitioners have increasingly sought to generate images featuring personal styles [51, 61, 44, 30]. Fine-tuning methods [45, 51, 30] have been particularly prominent in this area. StyleDrop [51], a study closely related to our work, utilizes a masked generative model [3] and involves human data selection through multi-stage training. Some works focus on learning multiple concepts simultaneously [27, 20] or merging several fine-tuned models [49, 30], while others analyze the diffusion model’s U-Net [43] layers to identify those most effective for learning styles  [9]. However, fine-tuning requires hyperparameter searches with each new large-scale model release [38, 8, 29], and is often applied without a deep understanding of the diffusion objectives.
As an alternative approach for style-driven generation, zero-shot approaches have been proposed [61, 44, 14, 59], but these methods still fall short in style alignment compared to fine-tuning and are often limited to specific domains [13, 46]. Due to these limitations, we focus on fine-tuning approaches in our work, aiming to provide insights into the behavior of diffusion objectives to make fine-tuning more accessible and effective.
@@ 6 Conclusion
In this paper, we observed that stylistic features in diffusion models emerge predominantly at higher noise levels. To address the limitations of previous fine-tuning approaches in capturing new artistic styles, we proposed the Style-friendly SNR sampler, which biases the SNR distribution towards higher noise levels. We showed style-driven generation that reflects reference styles in terms of color schemes, layouts, illumination, and brushstrokes. We hope this work will serve as a stepping stone toward using diffusion models as digital art previewers.
@@ Acknowledgement
Appendix
Appendix A Additional Results
A.1 Qualitative Results
Additional Samples.
Additional Comparison.
Effect of Varying $\mu$ on Object References.
We examine how the parameter $\mu$ in our Style-friendly SNR sampler impacts object reference fine-tuning in Fig. S5. While FLUX fine-tunes object references well using the SD3 sampler, setting $\mu=0$ leads to failures in color binding and structure. For example, generating a clock results in missing color and numeral; sunflowers around a backpack and vase cause the objects to turn yellow; and a purple rug makes the object also appear purple. In the flat cartoon devil example, the model fails to capture the reference’s unique short limbs and structure. These findings indicate that a larger $\mu$ hinders the learning of stylistic elements including color schemes inherent to objects. This suggests that the success of FLUX in object fine-tuning arises from its noise level distribution being adjusted toward object-centric generation.
SD3.5 Samples.
We extend our qualitative comparison by evaluating our Style-friendly SNR sampler using the SD3.5-8B model [57], comparing it against previous fine-tuning methods, namely the SD3 sampler [8] and DCO [30]. As shown in Fig. S6, the results are consistent with the qualitative comparisons using FLUX-dev presented in Fig. 7.
A.2 Quantitative Results
CLIP Scores.
In Fig. 6, we presented analyses of the mean $\mu$, standard deviation $\sigma$, and LoRA rank using the DINO similarity score. In Fig. 7(a), we provide the corresponding CLIP image similarity (CLIP-I) scores to further validate our findings. The CLIP-I scores exhibit a similar trend to the DINO scores, where decreasing $\mu$ enhances style alignment. Varying $\sigma$ affects the CLIP-I scores consistently with the DINO results. Our Style-friendly SNR sampler with $\mu=-6$ and a rank of 4 still outperforms the SD3 sampler with a rank of 32 (dotted lines).
Effectiveness Compared to Increasing Model Capacity.
To demonstrate that our method is more effective than increasing model capacity, we conduct an additional experiment where we fine-tune the model using the SD3 sampler with a higher LoRA rank of 128. As shown in Tab. S1, our Style-friendly SNR sampler with a rank of 32 achieves higher DINO and CLIP-I scores compared to the SD3 sampler with a rank of 128. This indicates that focusing on the critical noise levels where styles emerge has a more significant impact than increasing the number of trainable parameters. Therefore, our approach enhances style alignment without necessitating larger model capacities.
Trainable Parameters.
Figure S5: Varying $\mu$ on object references. The object names are written at the top of the reference images. Setting $\mu=0$ (high log-SNR value) leads to failures in color binding and structure when fine-tuning on object references, whereas using the SD3 sampler allows FLUX to fine-tune object references effectively. This unveils why recent diffusion models perform well on object fine-tuning, as their noise level distributions are adjusted toward object-centric generation.
Figure S6: Comparison of fine-tuning the SD3.5-8B. The results with SD3.5-8B are consistent with the qualitative comparison based on FLUX-dev presented in Fig. 7.
Method Model Metrics
DINO $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$
SD3 Sampler [8] FLUX-dev 0.373 0.645 0.350
      w/ rank 128 FLUX-dev 0.426 0.668 0.345
Style-friendly FLUX-dev 0.461 0.686 0.344
Table S1: Comparison to increasing LoRA rank.
Method DINO CLIP-I CLIP-T
Style-friendly 0.489 0.698 0.349
w/o Text attn 0.462 0.693 0.349
Table S2: Ablation study on trainable parameters.
(a) Varying $\mu$.
(b) Varying $\sigma$.
(c) Varying LoRA Rank.
A.3 Additional Observations
Styles Emerge at Higher Noise Levels.
We examine the case where style descriptions are given only during the initial denoising steps. Specifically, we condition the diffusion model with the prompt $y_{\text{w/ style}}$ for the first 10% of the denoising process and then switch to the non-style prompt $y_{\text{w/o style}}$ for the remaining steps. As shown in Fig. S8, the generated images still reflect the intended styles despite the absence of style descriptions during most of the denoising. This confirms that stylistic features emerge primarily at higher noise levels, corresponding to the early denoising steps. The images are similar to those generated when style descriptions are used throughout all steps, indicating that early conditioning suffices for style representation.
Probability Distribution of Timestep.
We present the probability density plot of timesteps corresponding to sampled log-SNR values. By negatively shifting the log-SNR distribution, the resulting timestep distribution becomes aggressively skewed toward $t=1$, as shown in Fig. S9. The colored region indicates the timestep interval where we observe styles emerging. This shift increases the probability density in the timesteps critical for style representation, focusing on these timesteps during fine-tuning for style-driven generation.
Figure S8: Prompt switching during generation. The generated images still reflect the intended styles even without style descriptions in most of the denoising process, indicating that stylistic features emerge mainly at the early denoising steps.
Figure S9: Probability distribution of timestep $t$. Colored region indicates style-emerging noise levels. The timestep $t$ ranges between 0 and 1, where $t=1$ corresponds to pure noise. As $t$ increases, the log-SNR $\lambda_{t}$ decreases.
Appendix B Baselines
In this section, we review the baseline methods used for comparison and highlight how our approach differs from them in addressing style-driven generation. The baseline methods include the fine-tuning-based methods—SD3 sampler [8], Direct Consistency Optimization (DCO) [30], and our method—as well as the zero-shot methods IP-Adapter [61], RB-Modulation [44], and Style-Aligned [14]. All fine-tuning-based methods are trained using the same reference images and training prompts to ensure a fair comparison. Notably, our Style-friendly SNR sampler disables timestep shifting [8] during fine-tuning. For the fine-tuning process, we utilize LoRA [19] with a rank of 32 (except the LoRA rank analysis) and a learning rate of $10^{-4}$. The models are trained with a batch size of 1, employing gradient accumulation over 4 steps. FLUX-dev is a guidance-distilled model [34] that provides the guidance scale as an input; therefore, during fine-tuning, we set the guidance scale to 1 to match standard diffusion fine-tuning. All experiments are implemented using the Hugging Face Diffusers library version 0.31.0 and conducted on a single NVIDIA A40 GPU. The FLUX-dev model is trained with gradient checkpointing to save memory consumption during training. All fine-tuning and inference using FLUX-dev are conducted with the BF16 quantization, while SD3.5-8B uses FP16.
B.1 Direct Consistency Optimization
$\mathcal{L}_{DCO}(x_{0})=\mathbb{E}_{t,\epsilon}\bigg{[}-\text{log}\sigma(-\beta Tw(t)$
$||f_{\theta}(x_{t},t)-f(x_{t},t)||^{2}-||f_{\phi}(x_{t},t)-f(x_{t},t)||^{2})\bigg{]}.$ (8)
B.2 IP-Adapter
B.3 RB-Modulation
RB-Modulation [44] is a zero-shot approach using Stable Cascade [37], a model accepting both CLIP image embeddings and text embeddings as inputs. During the denoising process, RB-Modulation employs gradient guidance of a CSD, a model fine-tuned from CLIP to measure style similarity, resembling classifier guidance [5]. At each denoising step, CSD computes the similarity between the approximated $x_{0}$ and the reference image, guiding the generation process to enhance this similarity. RB-Modulation also aggregates multiple attention features.
B.4 Style-Aligned
B.5 Offset Noise
Method Model Metrics
DINO $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$
SD3 Sampler [8] SD3.5 0.424 0.670 0.350
      w/ offset 0.1 SD3.5 0.452 0.678 0.353
Style-friendly SD3.5 0.489 0.698 0.349
      w/ offset 0.01 SD3.5 0.476 0.697 0.350
SD3 Sampler [8] FLUX-dev 0.373 0.645 0.350
      w/ offset 0.1 FLUX-dev 0.451 0.679 0.349
Style-friendly FLUX-dev 0.461 0.686 0.344
      w/ offset 0.01 FLUX-dev 0.500 0.704 0.341
Table S3: Incorporating offset noise. Offset noise improves SD3 sampler but still does not reach the performance of our Style-friendly SNR sampler; combining our Style-friendly approach with Offset Noise at a smaller scale (0.01) slightly enhances the style alignment of FLUX-dev.
Appendix C Experimental Details
C.1 Style Prompts
We conduct all quantitative evaluations using the 18 reference styles shown in the appendix of the StyleDrop paper [51]. The style prompts for these 18 styles can also be found in the StyleDrop appendix. For qualitative evaluations, we use additional challenging style references and we display the corresponding style prompt for each image in Fig. S10. The first and second references are produced with open-source meme generator source1 and source2 respectively. The third reference is cropped from the appendix of Dreambooth [45] paper.
Figure S10: Style prompts used for experiments.
C.2 Evaluation Prompts
@@
C.3 User Study

### v3

@@ 4.1 Analysis of Style-Friendly SNR Sampler
In Fig. 5 and Fig. 6, we conduct experiments to analyze the impact of varying the parameters of our Style-friendly SNR sampler—specifically, the mean ($\mu$) and the standard deviation ($\sigma$) of the log-SNR sampling distribution, as well as the LoRA rank.
@@
We experiment with $\mu$ values ranging from 0 to $-8$ for both FLUX-dev and SD3.5-8B. As shown in Fig. 5, increasing $\mu$ progressively impairs the model’s ability to capture reference styles. This trend is quantitatively confirmed in Fig. 6(a), where DINO similarity scores decrease with increasing $\mu$ (towards zero), clearly indicating poorer style alignment. Conversely, in Fig. 5, when $\mu$ is set to $-6$, the models begin to capture and reflect the reference styles effectively.
@@
We also investigate the effect of varying the standard deviation $\sigma$ of the log-SNR sampling distribution in Fig. 5 and Fig. 6(b). When $\sigma$ is less than 2, the model does not sufficiently cover the style-emerging region as shown by green curve of Fig. 2, leading to lower style alignment due to limited exposure to critical noise levels. In Fig. 5, $\sigma\geq 2$ reflects the ‘glowing’ style. Also, in Fig. 6(b), style similarity remains low for $\sigma<2$ but improves when $\sigma\geq 2$. For FLUX, DINO score is highest at $\sigma=3$, whereas for SD3.5, $\sigma=2$ yields the best results. Based on these findings, we adopt these hyperparameters for each respective model.
@@ 4.2 Qualitative Results
In Fig. 7, our Style-friendly SNR sampler accurately captures the styles of reference images, reflecting stylistic features including color schemes, layouts, illumination, and brushstrokes. In contrast, fine-tuning FLUX-dev with the standard SD3 sampler often fails to capture key stylistic components, such as layout (column 1), color scheme (columns 2-4), and illumination (column 5).
Fine-tuning FLUX-dev with DCO struggles to learn the reference styles due to strong regularization that prevents significant deviation from the pre-trained model. As seen in column 4 of the IP-Adapter results, woman in the reference appears, indicating content leakage. IP-Adapter with FLUX-dev and RB-Modulation rely on embeddings of CLIP [42] and CSD [54], which may not capture fine stylistic details, leading to less accurate style reproduction. Style-Aligned shares self-attention features within the diffusion model, which can cause artifacts such as destroyed structure (columns 1-3) when attention features conflict. While GPT-4o prompt utilizes the detailed style description, it often fails to reflect key stylistic features such as color scheme, highlighting the necessity of image guidance for effective style-driven generation.
4.3 Quantitative Results
@@
We further conduct a user study to quantify human preferences using Amazon Mechanical Turk. Following previous work [53], we compare our method to each method with two separate questionnaires. According to the reference style image and target text prompt, users are asked to select which of the two generated images is more similar to the style in the reference image and represents the target text prompt better. We obtain 450 answers from 150 participants for each comparison, and the results are presented in Tab. 1. Our method outperforms prior arts in both aspects ($p<0.05$ in the Wilcoxon signed-rank test), consistent with the qualitative results and demonstrates the superiority of our method in learning stylistic elements. More details on our user study are provided in the Appendix.
In Tab. 2, we evaluate our method and prior works using DINO [3] and CLIP image similarities (CLIP-I) to assess style alignment, and CLIP text-image similarity (CLIP-T) for text alignment. Our method achieves the highest DINO and CLIP-I scores, demonstrating its superior ability to capture styles from the reference images. Notably, our approach improves DCO, indicating that the key factor in style-driven fine-tuning is not whether the method relies on diffusion loss or reinforcement learning-based loss, but rather which noise levels are emphasized during training.
While our CLIP-T is slightly lower compared to some methods, we already showed superior text alignment in human evaluation (Tab. 1). This discrepancy arises because textual style descriptions alone can be inherently ambiguous and often correspond to common interpretations. Consequently, methods that accurately reflect unique styles may deviate from these typical interpretations, leading to lower CLIP-T scores despite higher alignment to the intended reference styles. Overall, our quantitative results confirm that our method accurately reflects both styles and texts.
@@
IP-Adapter [62] FLUX 0.361 0.656 0.354
GPT-4o Prompt FLUX 0.299 0.621 0.338
StyleDrop [53] MUSE 0.465 0.665 0.325
SD3 sampler [9] SD3.5 0.424 0.670 0.350
+Style-friendly SD3.5 0.489 0.698 0.349
@@
+Style-friendly SD3.5 0.478 0.695 0.351
SD3 sampler [9] FLUX 0.373 0.645 0.350
+Style-friendly FLUX 0.478 0.691 0.343
DCO [32] FLUX 0.373 0.643 0.353
+Style-friendly FLUX 0.488 0.698 0.341
Table 2: Quantitative comparison. Style alignment (DINO and CLIP-I) and text alignment (CLIP-T) with 18 styles from [53]. Our style-friendly exhibits superior style-alignment scores. Rows 1-3 show non-fine-tuning baselines. †: Results obtained using an unofficial implementation [1].
@@ 4.4 Applications
While Dreambooth [47] paper demonstrates generating multi-panel comics by generating each panel with a fine-tuned diffusion model, we define the entire multi-panel comic itself as a unique style. Our method treats multiple panels as a single image during fine-tuning, enabling the generation of coherent multi-panel comics from only a single reference (see the first row of Fig. 8). By specifying a new subject in the target prompt, the model consistently places that subject across all comic-style panels. Beyond comics, our method also extends to typography, leveraging the spelling capabilities of recent models [31, 59]. As shown in the second row of Fig. 8, this flexibility allows users to effortlessly generate a broad range of customized textual elements in unique styles.
@@ 5.1 Diffusion Models
Figure 8: Multi-panel and typography. First row demonstrates generating multiple coherent panels as a single image. Second row shows customized typography with a unique style.
@@ 5.2 Style-Driven Generation
With advancements in text-to-image models, practitioners have increasingly sought to generate images featuring personal styles [53, 62, 46, 32]. Fine-tuning methods [47, 53, 32] have been particularly prominent in this area. StyleDrop [53], a study closely related to our work, utilizes a masked generative model [4] and involves human data selection through multi-stage training. Some works focus on learning multiple concepts simultaneously [29, 21] or merging several fine-tuned models [51, 32], while others analyze the diffusion model’s U-Net [45] layers to identify those most effective for learning styles  [10]. However, several existing methods have not yet been validated or applied to recently released large-scale models [9, 31]. In contrast, our method provides an adaptable fine-tuning strategy, which we validate on up-to-date diffusion models.
As an alternative approach for style-driven generation, zero-shot approaches have been proposed [62, 46, 15, 60], but these methods still fall short in style alignment compared to fine-tuning and are often limited to specific domains [14, 48]. Furthermore, some methods [46, 15] require extra inference-time gradient guidance [46] or inversion [15], increasing inference cost. Due to these limitations, we focus on fine-tuning in our work, aiming to provide insights into the behavior of diffusion objectives to make fine-tuning more accessible and effective. While fine-tuning is not the only option, our results show that it is a promising approach.
@@ 6 Conclusion
In this paper, we observed that stylistic features in diffusion models emerge predominantly at higher noise levels. To address the limitations of previous fine-tuning approaches in capturing new artistic styles, we proposed the Style-friendly SNR sampler, which biases the SNR distribution towards higher noise levels. We showed style-driven generation that reflects reference styles specified via text prompts. We hope this work will serve as a stepping stone toward using diffusion models as digital art previewers.
@@ Acknowledgement
Supplementary Material
Appendix A Experimental Details
A.1 Style Prompts
We conduct all quantitative evaluations using the 18 reference styles shown in the appendix of the StyleDrop paper [53]. The style prompts for these 18 styles can also be found in the StyleDrop appendix.
A.2 Evaluation Prompts
@@
In Fig. S1, we present the detailed style descriptions generated by GPT-4o, which serve as text prompts for the GPT-4o Prompt baseline. Specifically, for each of the 18 reference styles, we obtained comprehensive textual descriptions using GPT-4o. These detailed prompts were directly used for text-to-image generation.
A.3 User Study
@@ A.4 Implementation
Optimizer and Learning Rate. We use the Adam optimizer [28] at a learning rate of 10-4 for 300 steps. The batch size is set to 1, with gradient accumulation over 4 steps.
Guidance Scale and Inference Steps. During inference, we use a guidance scale [17, 37] of 7.0. The number of denoising steps is set to 28.
Model-Specific Details. FLUX-dev [31] is a guidance-distilled model [37] that takes guidance scale as input. For fine-tuning, we fix the guidance scale to 1.0 to match standard diffusion training, enable gradient checkpointing for memory efficiency, and use BF16 quantization for both fine-tuning and inference. SD3.5-8B [59] use FP16 precision.
Disabling Timestep Shifting. SD3 [9] uses a timestep shifting mechanism. However, we disable this shifting for our Style-friendly SNR sampler to isolate the effect of our proposed SNR sampling strategy.
Baseline Implementation. We use the Hugging Face Diffusers library (version 0.31.0) for consistent training and inference across methods.
Figure S1: Detailed style description. These descriptions serve as the input prompts for our GPT-4o [23] prompt baseline, which generates images solely based on these textual style specifications.
@@
Appendix B Baselines
B.1 Direct Consistency Optimization
$\mathcal{L}_{DCO}(x_{0})=\mathbb{E}_{t,\epsilon}\bigg{[}-\text{log}\sigma(-\beta T$
$||v_{\theta}(x_{t},t)-v(x_{t},t)||^{2}-||v_{\phi}(x_{t},t)-v(x_{t},t)||^{2})\bigg{]},$ (7)
where $v(x_{t},t)$ is target velocity field, $v_{\theta}$, is fine-tuning model, and $v_{\phi}$ is frozen pre-trained model.
B.2 IP-Adapter
B.3 RB-Modulation
RB-Modulation [46] is a zero-shot approach using Stable Cascade [40], a model accepting both CLIP image embeddings and text embeddings as inputs. During the denoising process, RB-Modulation employs gradient guidance of a CSD [54], a model fine-tuned from CLIP to measure style similarity, resembling classifier guidance [7]. At each denoising step, CSD computes the similarity between the approximated $x_{0}$ and the reference image, guiding the generation process to enhance this similarity. RB-Modulation also aggregates multiple attention features.
B.4 Style-Aligned
B.5 Offset Noise
Method Model Metrics
DINO $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$
SD3 Sampler [9] SD3.5 0.424 0.670 0.350
      w/ offset 0.1 SD3.5 0.452 0.678 0.353
Style-friendly SD3.5 0.489 0.698 0.349
      w/ offset 0.01 SD3.5 0.476 0.697 0.350
SD3 Sampler [9] FLUX-dev 0.373 0.645 0.350
      w/ offset 0.1 FLUX-dev 0.451 0.679 0.349
Style-friendly FLUX-dev 0.461 0.686 0.344
      w/ offset 0.01 FLUX-dev 0.500 0.704 0.341
Table S1: Incorporating offset noise. Offset noise improves SD3 sampler but still does not reach the performance of our Style-friendly SNR sampler; combining our Style-friendly approach with Offset Noise at a smaller scale (0.01) slightly enhances the style alignment of FLUX-dev. Here, we use $\sigma=2$ for Style-friendly.
Appendix C Additional Results
C.1 Quantitative Results
(a) Varying $\mu$.
(b) Varying $\sigma$.
(c) Varying LoRA Rank.
CLIP Scores.
In the main paper, we presented analyses of the mean $\mu$, standard deviation $\sigma$, and LoRA rank using the DINO similarity score. In Fig. 4(a), we provide the corresponding CLIP image similarity (CLIP-I) scores to further validate our findings. The CLIP-I scores exhibit a similar trend to the DINO scores, where decreasing $\mu$ enhances style alignment. Varying $\sigma$ affects the CLIP-I scores consistently with the DINO results. Our Style-friendly SNR sampler with $\mu=-6$ and a rank of 4 still outperforms the SD3 sampler with a rank of 32 (dotted lines).
Method Model Metrics
DINO $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$
SD3 Sampler [9] FLUX-dev 0.373 0.645 0.350
      w/ rank 128 FLUX-dev 0.426 0.668 0.345
Style-friendly FLUX-dev 0.461 0.686 0.344
Table S2: Comparison to increasing LoRA rank.
Method DINO CLIP-I CLIP-T
Style-friendly 0.489 0.698 0.349
w/o Text attn 0.462 0.693 0.349
Table S3: Ablation study on trainable parameters.
Effectiveness Compared to Increasing Model Capacity.
To demonstrate that our method is more effective than increasing model capacity, we conduct an additional experiment where we fine-tune the model using the SD3 sampler with a higher LoRA rank of 128. As shown in Tab. S2, our Style-friendly SNR sampler with a rank of 32 achieves higher DINO and CLIP-I scores compared to the SD3 sampler with a rank of 128. This indicates that focusing on the critical noise levels where styles emerge has a more significant impact than increasing the number of trainable parameters.
Trainable Parameters.
C.2 Qualitative Results
SD3.5 Samples.
We extend our qualitative comparison by evaluating our Style-friendly SNR sampler using the SD3.5-8B model [59], comparing it against previous fine-tuning methods, namely the SD3 sampler [9] and DCO [32]. As shown in Fig. S5, the results are consistent with the qualitative comparisons using FLUX-dev presented in the main paper.
Additional Comparison.
Object fine-tuning using Style-friendly SNR sampler.
We further evaluate whether our proposed Style-friendly SNR sampler—designed specifically for style learning—affects object-driven generation performance. As shown in Fig. S8, our Style-friendly SNR sampler and the original SNR sampler produce qualitatively similar results when fine-tuning on object references. Our approach successfully captures critical details, including shapes, colors, and prominent text or patterns (e.g., on the bowl and can). However, it occasionally omits subtle features, such as the small teeth of the monster toy. These findings support our hypothesis that distinct approaches are required for object-centric and style-driven fine-tuning; our Style-friendly sampler, while slightly suboptimal for object-centric fine-tuning, excels in capturing nuanced style characteristics.
Additional Samples.
Figure S5: Comparison of fine-tuning the SD3.5-8B. We show ‘A crown’, ‘A Golden Gate bridge’, and ‘An espresso machine’ in various styles. The results with SD3.5-8B are consistent with the qualitative comparison based on FLUX-dev presented in the main paper.
Figure S8: Object fine-tuning comparison. We compare our Style-friendly SNR sampler and the standard sampler on object-driven fine-tuning. Both approaches generate similar overall results, though our Style-friendly sampler occasionally misses minor details, such as the small teeth of the monster toy. Nevertheless, the Style-friendly sampler reliably captures the object’s overall shape, color, and key details such as text and patterns on the bowl and can. The object names are written at the top of the reference images.